OpenStudy (anonymous):

Simplify the following quotient of complex numbers into the form a + bi. (-4-3i)/(2-i)

OpenStudy (anonymous):

multiply top and bottom by conjugate of the denominator conjugate of \(a+bi\) is \(a-bi\) and this works because \((a+bi)(a-bi)=a^2+b^2\) a real number

OpenStudy (anonymous):

in your case you will get \[\frac{-4-3i}{2-i}\times \frac{2+i}{2+i}=\frac{-(4-3i)(2+i)}{2^2+1^2}\]

OpenStudy (anonymous):

typo there \[\frac{-4-3i}{2-i}\times \frac{2+i}{2+i}=\frac{(-4-3i)(2+i)}{2^2+1^2}\]

OpenStudy (anonymous):

denominator is 5, numerator is what you get when you multiply \((-4-3i)(2+i)\) and btw this should be called "divide" not "simplify" just thought i would mention it

OpenStudy (anonymous):

so the answer would be -11-10x/5 right

OpenStudy (unklerhaukus):

the other way is to convert to the numerator and denominator \(re^{i\theta}\) then division is simple